# Learning Motion Blur Robust Vision Transformers with Dynamic Early Exit for Real-Time UAV Tracking

### v1

@@ IV-E Comparison with Deep Trackers
The proposed BDTrack-DeiT is also compared with 12 SOTA deep trackers, i.e., PrDiMP50 [79], AutoMatch [77], SOAT [78], CSWinTT [76], SimTrack[14], OSTrack [13], ZoomTrack [15], SeqTrack [70], MAT [74], SuperSBT [71], DCPT [73], and EVPTrack [72]. We utilize Table II to present the results of precision (Prec.) and success rate (Succ.), their average (Avg.), and average GPU Speed (Avg.FPS). Our BDTrack-DeiT stands out by offering the fastest speed without compromising on high performance, showcasing its competitive edge in terms of both performance and speed. Remarkably, our method achieves second place in both Avg. precision on five benchmarks and precision on UAVDT, while securing the top position in precision on VisDrone2018. This underscores the effectiveness of our approach for UAV tracking tasks and highlights that our BDTrack-DeiT is equivalent in precision to state-of-the-art deep trackers. Although other deep trackers like SeqTrack [70] and EVPTrack [72] may achieve comparable performance with ours, their GPU speeds are significantly slower. In fact, our method outperforms SeqTrack and EVPTrack by being 9 and 11 times faster, respectively. By showcasing its capability to offer high precision and speed, these results demonstrate the suitability of our approach for real-time UAV tracking applications that prioritize accuracy and efficiency.
@@
TABLE II: Precision (Prec.), success rate (Succ.), and GPU speed comparison between BDTrack-DeiT and DL-based trackers.
DTB70 UAVDT VisDrone2018 UAV123 UAV123@10fps Avg.
Tracker Source Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Avg.FPS
PrDiMP50[79] CVPR 20 84.0 64.3 75.8 55.9 79.8 60.2 87.2 66.5 83.9 64.7 82.1 62.3 53.6
AutoMatch[77] ICCV 21 82.5 63.4 82.1 60.8 78.1 59.6 83.8 64.4 78.1 59.4 80.9 61.9 63.9
SOAT[78] ICCV 21 83.1 64.6 82.1 60.7 76.9 59.1 82.7 64.9 85.2 65.7 82.0 63.0 36.2
CSWinTT[76] CVPR 22 80.3 62.3 67.3 54.0 75.2 58.0 87.6 70.5 87.1 68.1 79.5 62.6 12.3
SimTrack[14] ECCV 22 83.2 64.6 76.5 57.2 80.0 60.9 88.2 69.2 87.5 69.0 83.1 64.2 72.8
OSTrack[13] ECCV 22 82.7 65.0 85.1 63.4 84.2 64.8 84.7 67.4 83.1 66.1 83.9 65.3 68.4
ZoomTrack[15] NIPS 23 82.0 63.2 77.1 57.9 81.4 63.6 88.4 69.6 88.8 70.0 83.5 64.8 62.7
SeqTrack[70] CVPR 23 85.6 65.5 78.7 58.8 83.3 64.1 86.8 68.6 85.7 68.1 84.0 65.0 32.3
MAT[74] CVPR 23 83.2 64.5 72.9 54.8 81.6 62.2 86.7 68.3 86.9 68.5 82.3 63.6 71.2
SuperSBT[71] arXiv’24 84.4 65.1 80.1 59.2 82.2 63.5 85.3 67.3 87.3 69.1 83.9 64.8 32.5
DCPT[73] ICRA 24 84.0 64.8 76.8 56.9 83.1 64.2 85.7 68.1 86.9 69.1 83.3 64.6 39.2
EVPTrack[72] AAAI 24 85.8 66.5 80.6 61.2 84.5 65.8 88.9 70.2 88.7 70.4 85.7 66.8 26.1
BDTrack-DeiT Ours 83.5 64.1 84.1 61.0 85.2 64.3 84.8 66.7 83.5 65.9 84.2 64.4 287.2
TABLE III: Comparison of FLOPS, parameters, and inference speed on NVIDIA Jetson AGX Xavier edge device (AGX.FPS) with ten efficient lightweight trackers, evaluated on UAVDT.
Tracker Source FLOPs Params. AGX.FPS
BDTrack-DeiT 1.7G-2.4G 5.8M-8.0M 44.9
BDTrack-ViT 1.7G-2.4G 5.8M-8.0M 40.9
BDTrack-T2T Ours 1.9G-2.4G 5.4M-7.0M 40.1
LightFC [54] KBS 24 2.9G 7.5M 34.7
SMAT [53] WACV 24 3.2G 8.6M 33.8
Aba-ViTrack [18] ICCV 23 2.4G 8.0M 37.3
HiT [51] ICCV 23 1.3G 11.0M 41.2
ABDNet [33] RAL 23 8.27G 12.3M 33.2
SGDViT [56] ICRA 23 11.3G 23.3M 32.1
DRCI [57] ICME 23 3.6G 8.8M 44.6
TCTrack [5] CVPR 22 6.9G 8.5M 34.4
HCAT [58] ECCV 22 2.6G 8.4M 34.8
HiFT [1] ICCV 21 7.2G 9.9M 35.6
TABLE IV: Ablation study of the impact of MBRV and DEEM on the baseline trackers’ performance on DTB70, UAVDT, VisDrone2018, UAV123, and UAV123@10fps. Note that precision and success rate are shown in the form of (Prec. , Succ.)
Tracker MBRV DEEM DTB70 UAVDT VisDrone2018 UAV123 UAV123@10fps Avg. Avg.FPS
(79.3, 62.4) (77.0, 55.6) (83.0, 62.7) (83.2, 66.5) (82.1, 64.8) (80.9, 62.4) 195.4
$\checkmark$ (83.5, 64.7) (79.3, 57.4) (84.6, 63.8) (84.7, 66.9) (84.5, 66.4) (83.3↑2.4, 63.8↑1.4) -
BDTrack-ViT $\checkmark$ $\checkmark$ (82.8, 64.2) (78.9, 57.3) (83.9, 63.6) (83.5, 65.5) (84.1, 66.1) (82.6↑1.7, 63.3↑0.9) 238.6↑22%
(78.6, 62.0) (73.5, 53.7) (81.2, 62.1) (81.9, 64.9) (79.2. 63.2) (78.9, 61.2) 202.1
$\checkmark$ (81.6, 63.6) (78.5, 56.8) (82.8, 63.1) (81.5, 64.3) (83.3, 65.3) (81.5↑2.6, 62.6↑1.4) -
DBTrack-T2T $\checkmark$ $\checkmark$ (80.7, 63.0) (77.9, 56.3) (81.4, 62.2) (80.4, 63.4) (82.9, 65.1) (80.6, 62.0) 240.4↑19%
(84.2, 65.1) (78.6, 56.7) (81.6, 62.2) (83.7, 66.1) (82.7, 65.6) (82.2, 63.2) 232.5
$\checkmark$ (84.9, 65.6) (84.5, 61.4) (86.1, 65.1) (85.3, 67.0) (84.2, 66.2) (84.8↑2.6, 65.0↑1.8) -
DBTrack-DeiT $\checkmark$ $\checkmark$ (83.5, 64.1) (84.1, 61.0) (85.2, 64.3) (84.8, 66.7) (83.5, 65.9) (84.2↑2.0, 64.4↑1.2) 287.2↑24%
IV-F Efficiency Comparison
To highlight the superiority of the proposed trackers over existing ones in terms of efficiency, we present in Table III the floating-point operations per second (FLOPs), the number of parameters during inference (Param.), and the speed on the Nvidia Jetson AGX Xavier edge device (AGX.FPS) of our method and 10 state-of-the-art efficient lightweight trackers. Notably, as our methods feature adaptive architectures, the FLOPs and Params. of our method vary within a range, spanning from the minimum to the maximum values. As observed, our methods achieve significantly lower FLOPs and Params. compared to all rival trackers, except for HiT, which secures the lowest FLOPs. Even our maximum values are lower than those of most trackers. Although HiT has the lowest FLOPs, its Params. is significantly larger than ours, and its tracking performance across multiple UAV tracking benchmarks is notably inferior, as shown in Table I. In terms of AGX speed, our BDTrack-DeiT attains the highest speed of 44.9 FPS. Additionally, our BDTrack-ViT and BDTrack-T2T, with speeds exceeding 40 FPS, also surpass most rival trackers.
IV-G Ablation Study
@@
IV-G1 Impact of learning blur robust ViTs with dynamic early exit
@@
We add the proposed components MBRV and DEEM into the baselines one by one to evaluate their effectiveness. The evaluation results on five benchmarks are shown in Table IV. As GPU speed of the baseline and its enhanced version by MBRV are theoretically equal, we provide only the speed of the baseline to eliminate potential nuances arising from randomness. As can be seen, the inclusion of MBRV leads to significant improvements in both Prec. and Succ. for all baseline trackers. Specifically, the inclusion of the MBRV leads to a significant increase of over 2.0% in average Prec. and 1.0% in average Succ. for all baseline trackers. When DEEM is further integrated, consistent improvements in GPU speeds are observed, accompanied by only slight decreases in Prec. and Succ. Specifically, the GPU speeds of all baselines see increases of above 19.0%, with BDTrack-DeiT achieving a remarkable 24.0% improvement. Despite the marginal drop in Prec. and Succ. (the average drops are less than 0.9%), the overall pattern of enhanced GPU speeds and minor sacrifices in tracking performance validates the effectiveness of DEEM in optimizing tracking efficiency.
Figure 6: Ablation study on when to initiate the DEEM. The figure presents the results for average precision (Avg. Prec.), average success rate (Avg. Succ.), and average tracking GPU speed (Avg. FPS) with respect to $n_{enf}$ across the five datasets.
IV-G2 Study on when to initiate the DEEM
We conducted experiments to examine the impact of the timing of initiating the DEEM on performance by varying the parameter $n_{enf}$ from 3 to 8. Fig. 6 presents the results of average precision (Avg. Prec.), average success rate (Avg. Succ.), and average GPU speed (Avg. FPS) across the five datasets. As shown, the timing of initializing the DEEM in BDTrack-DeiT directly impacts both the tracking performance and speed. Specifically, increasing $n_{enf}$ results in an upward trend for both Avg. Prec. and Avg. Succ., but a decrease in Avg. FPS. When $n_{enf}$ falls in the range of 3 to 5, each additional layer leads to about 2.0% increase in both Avg. Prec. and Avg. Succ., while causing about 10% drop in speed. When the value of $n_{enf}$ exceeds 6, further increases in $n_{enf}$ do not result in significant improvements in precision and success rate, while the speed continues to gradually decrease significantly. Considering the trade-off between performance and speed, we have chosen to set the default value of $n_{enf}$ as 6 in our implementation.
@@ IV-G2 Study on when to initiate the DEEM
TABLE V: Ablation study on the weighting of loss $\mathcal{L}_{br}$ in training BDTrack-DeiT by varying $\rho$ from $0.5\times 10^{-4}$ to $5.0\times 10^{-4}$. Note that precision and success rate are shown in the form of (Prec. , Succ.).
$\rho\times 10^{-4}$ DTB70 UAVDT VisDrone2018 UAV123 UAV123@10fps Avg.
0.5 (82.2, 63.6) (83.4, 60.0) (83.7, 63.5) (83.1, 65.4) (83.2, 65.6) (83.1, 63.6)
$0.6$ (82.9, 63.9) (82.3, 59.3) (84.5, 63.9) (83.7, 66.0) (83.0, 65.4) (83.3, 63.7)
$0.7$ (83.2, 63.8) (83.0, 59.6) (85.8, 64.6) (84.1, 66.5) (83.3, 66.2) (83.9, 64.1)
$0.8$ (84.5, 64.8) (83.2, 59.8) (81.1, 61.9) (83.5, 65.8) (82.3, 64.9) (82.9, 63.4)
$0.9$ (81.8, 62.9) (82.5, 59.7) (84.6, 63.9) (84.0, 66.1) (83.1, 65.6) (83.2, 63.6)
$1.0$ (83.5, 64.1) (84.1, 61.0) (85.2, 64.3) (84.8, 66.7) (83.5, 65.9) (84.2, 64.4)
$2.0$ (83.2, 64.5) (80.4, 58.2) (86.1, 64.8) (84.0, 65.9) (84.9, 66.5) (83.7, 64.0)
$3.0$ (83.4, 64.3) (83.8, 60.7) (85.6, 64.5) (82.4, 64.7) (80.2, 63.6) (83.1, 63.6)
$4.0$ (81.6, 62.8) (82.4, 59.4) (84.5, 63.8) (84.5, 66.5) (80.0, 63.5) (82.6, 63.2)
$5.0$ (84.1, 64.6) (82.6, 59.7) (83.9, 63.6) (84.0, 66.0) (82.3, 65.1) (83.4, 63.8)
TABLE VI: Ablation study on the weighting of loss $\mathcal{L}_{spar}$ in training BDTrack-DeiT by varying $\gamma$ from $0.5\times 10^{3}$ to $1.5\times 10^{3}$. Note that precision and success rate are shown in the form of (Prec. , Succ.).
$\gamma\times 10^{3}$ DTB70 UAVDT VisDrone2018 UAV123 UAV123@10fps Avg.
$0.5$ (84.1, 65.0) (81.2, 58.6) (83.7, 63.1) (83.4, 65.6) (81.5, 64.8) (82.8, 63.5)
$0.6$ (82.5, 64.4) (81.1, 58.7) (86.7, 65.3) ( 84.1, 66.0) (82.3, 65.2) (83.3, 63.9)
$0.7$ (82.9, 64.4) (81.9, 59.0) (83.9, 63.7) (82.9, 65.2) (83.4, 65.7) (83.0, 63.6)
$0.8$ ( 83.7, 64.9) (84.2, 60.4) (83.2, 63.0) (83.6, 65.6) (82.7, 65.2) (83.5, 63.8)
$0.9$ (82.7, 64.2) ( 84.0, 60.7) ( 86.2, 64.9) (82.9, 65.3) (83.0, 65.4) ( 83.8, 64.1)
$1.0$ (83.5, 64.1) (84.1, 61.0) (85.2, 64.3) (84.8, 66.7) ( 83.5, 65.9) (84.2, 64.4)
$1.1$ (81.6, 63.9) (83.5, 60.2) (84.7, 64.2) (83.4, 65.6) (83.0, 65.4) (83.3, 63.8)
$1.2$ (84.3, 65.1) (82.7, 59.8) (83.6, 63.2) (84.5, 66.2) (84.9, 66.5) (84.0, 64.2)
$1.3$ (83.1, 63.7) (82.7, 59.7) (84.6, 63.6) (83.8, 65.9) (84.1, 64.1) (83.7, 63.5)
$1.4$ (81.0, 61.9) (82.5, 59.9) (87.6, 66.1) (84.0, 65.9) (82.4, 63.5) (83.5, 63.4)
$1.5$ (81.2, 62.0) (83.9, 60.5) (83.7, 63.4) (82.7, 65.4) (81.7, 64.3) (82.7, 63.0)
@@
IV-G3 Study on weighting the loss for learning motion blur robust ViTs (MBRV)

### v2

@@ 4.3 Comparison with Deep Trackers
The proposed BDTrack-DeiT is also compared with 14 SOTA deep trackers. Table 2 presents the results in terms of precision (Prec.), success rate (Succ.), and their average (Avg.). It is worth noting that most deep trackers are tailored for generic tracking scenarios without prioritizing real-time constraints. Consequently, they tend to exhibit slower tracking speeds and higher model complexity, as shown in Table 3. Our BDTrack-DeiT stands out by offering the fastest speed without compromising on high performance, showcasing its competitive edge in terms of both performance and speed. Remarkably, our method achieves third place in Avg. precision on four benchmarks while securing the second position in precision on UAVDT and VisDrone2018. This underscores the effectiveness of our approach for UAV tracking tasks and highlights that our BDTrack-DeiT is equivalent in precision to state-of-the-art deep trackers. Although other deep trackers like ROMTrack [77] and EVPTrack [78] may achieve comparable performance with ours, their GPU speeds are significantly slower. In fact, our method outperforms SeqTrack and EVPTrack by being 5 and 11 times faster, respectively. By showcasing its capability to offer high precision and speed, these results demonstrate the suitability of our approach for real-time UAV tracking applications that prioritize accuracy and efficiency.
@@
Table 2: Precision (Prec.) and success rate (Succ.) comparison between BDTrack-DeiT and DL-based trackers.
UAVDT VisDrone2018 UAV123 UAV123@10fps Avg.
Tracker Source Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ.
PrDiMP [79] CVPR 20 75.8 55.9 79.8 60.2 87.2 66.5 83.9 64.7 81.7 61.8
SOAT [80] CVPR 21 82.1 60.7 76.9 59.7 82.7 64.9 85.2 65.7 81.7 62.8
AutoMatch [81] ICCV 21 82.1 62.9 78.1 59.6 83.8 64.4 78.1 59.4 80.5 61.6
SparseTT [82] IJCAI 22 82.8 65.4 81.4 62.1 85.4 68.8 82.2 64.9 82.9 65.3
CSWinTT [83] CVPR 22 67.3 54.0 75.2 58.0 87.6 70.5 87.1 68.1 79.3 62.7
SimTrack [27] ECCV 22 76.5 57.2 80.0 60.9 88.2 69.2 87.5 69.0 83.1 64.0
OSTrack [26] ECCV 22 85.1 63.4 84.2 64.8 84.7 67.4 83.1 66.1 84.2 65.2
ZoomTrack [84] NIPS 23 77.1 57.9 81.4 63.6 88.4 69.6 88.8 70.0 83.9 65.3
SeqTrack [85] CVPR 23 78.7 58.8 83.3 64.1 86.8 68.6 85.7 68.1 83.6 64.9
MAT [86] CVPR 23 72.9 54.8 81.6 62.2 86.7 68.3 86.9 68.5 82.0 63.5
ROMTrack [77] ICCV 23 81.9 61.6 86.4 66.7 87.4 69.2 85.0 67.8 85.2 66.3
DCPT [87] ICRA 24 76.8 56.9 83.1 64.2 85.7 68.1 86.9 69.1 83.1 64.5
EVPTrack [78] AAAI 24 80.6 61.2 84.5 65.8 88.9 70.2 88.7 70.4 85.7 66.9
SuperSBT [88] TPAMI 24 80.1 59.2 82.2 63.5 85.3 67.3 87.3 69.1 83.7 64.7
BDTrack-DeiT
Ours
84.1 61.0 85.2 64.3 84.8 66.7 83.5 65.9 84.4 64.5
Table 3: Comparison of FLOPS, parameters (Params.), and average speed between lightweight and deep trackers on four UAV tracking benchmarks.
Avg.FPS
Method FLOPs (G) Params. (M) GPU CPU
HiFT [7] 7.2 9.9 157.6 -
SiamAPN++ [17] 8.2 14.7 162.3 -
TCTrack [8] 16.9 8.5 133.8 -
SGDViT [67] 11.3 23.3 105.9 -
PRL-Track [69] 7.4 12.1 133.2 -
Aba-ViTrack [9] 2.4 7.9 177.8 49.7
LiteTrack [70] 28.3 7.3 115.6 -
AVTrack-DeiT [73] 0.97-2.4 3.5-7.9 253.6 58.8
Lightweight
BDTrack-DeiT(Ours)
1.7-2.4 5.8-7.9 283.4 63.2
OSTrack [26] 21.5 92.1 65.4 -
SeqTrack [85] 66.3 89.4 30.1 -
ZoomTrack [84] 21.2 91.8 60.8 -
MAT [86] 42.9 88.4 68.4 -
ROMTrack [77] 34.5 92.1 50.8 -
DCPT [87] 29.4 92.9 37.3 -
EVPTrack [78] 65.4 73.5 24.1 -
Deep SuperSBT [88] 24.6 65.5 30.3 -
4.4 Efficiency Comparison
To highlight the efficiency of the proposed trackers compared to existing methods, Table 3 presents the floating-point operations (FLOPs), the number of parameters during inference (Param.), and the inference speed of our method, alongside 8 SOTA efficient lightweight trackers and 8 SOTA deep trackers. Notably, as our method and AVTrack-DeiT feature adaptive architectures, the FLOPs and Params. of them vary within a range, spanning from the minimum to the maximum values. As observed, our methods achieve significantly lower FLOPs and Params. compared to all rival trackers. Even our maximum values are lower than those of most lightweight trackers. In terms of GPU speed, BDTrack-DeiT achieves the highest performance at 283.4 FPS, followed closely by AVTrack-DeiT [73] with 253.6 FPS. As for CPU speed, our tracker exhibits real-time performance on a single CPU [^1] Note that the real-time performance discussed in this paper can be generalized only to platforms similar to or more advanced than ours.. Although Aba-ViTrack [9] achieves comparable performance to BDTrack-DeiT, it exhibits significantly slower inference speed. Remarkably, BDTrack-DeiT boasts over 1.6 times the GPU speed and 1.3 times the CPU speed of Aba-ViTrack, striking a better balance between tracking precision and efficiency. On the other hand, although existing deep trackers achieve comparable performance to ours, they suffer from significantly slower speeds and higher model complexity. In contrast, our method outperforms ROMTrack [77] and EVPTrack [78] by factors of 5 and 11, respectively. These results further highlight the suitability of our approach for real-time UAV tracking applications where both accuracy and efficiency are critical.
Figure 4: The precision and success plots for attribute-based comparisons on ‘Motion Blur’ (top row) and ‘Camera Motion’ (bottom row) are presented for the attribute subsets of UAVDT. Note that BDTrack-DeiT* denotes BDTrack-DeiT without the application of the proposed components.
4.5 Attribute-Based Evaluation
To further validate the superiority of the proposed tracker over existing UAV trackers, we evaluate the performance of BDTrack-DeiT by comparing it with 14 SOTA UAV trackers on three subsets of the UAVDT [58], including ‘Motion blur’ and ‘Camera motion’. Note that we also evaluate BDTrack-DeiT without applying the components (i.e., MBRV and DEEM), denoted as BDTrack-DeiT* for reference. The precision and success plots are illustrated in Fig. 4. As observed, BDTrack-DeiT significantly outperforms the second tracker on ‘Motion blur’ and ‘Camera motion’, with gains of 1.2%/0.7% and 2.3%/1.8% in precision/success rate, respectively. Significantly, across all these two attributes, the incorporation of the proposed components yields a noteworthy enhancement over BDTrack-DeiT*, with improvements of 7.0%/5.0% and 7.9%/5.5%in terms of precision and success rate. Note that ‘Camera motion’ is the factor that can contribute to motion blur in visual tracking systems. When a camera moves rapidly or changes direction suddenly, it can cause motion blur in the captured images. Therefore, this attribute-based evaluation validates, directly or indirectly, the superiority and effectiveness of the proposed method in addressing the challenges posed by motion blur.
Figure 5: Visualization of qualitative comparisons between our tracker and six SOTA UAV trackers on three sequences from UAV123 [60], VisDrone2018 [59], and UAVDT [58] (i.e., person10, uav0000164_00000_s, and S1606).
4.6 Qualitative evaluation
Qualitative tracking results of BDTrack-DeiT and six SOTA trackers, i.e., HiFT [7], AutoTrack [24], TCTrack [8], PRL-Track [69], AVTrack-DeiT [73], and Aba-ViTrack [9], are presented in Fig. 5. Three video sequences from three different benchmarks are selected for demonstration: person10 from UAV123, uav0000164_00000_s from VisDrone2018, and S1606 from UAVDT. For better visualization, the low-resolution targets within the video frames are selectively zoomed, cropped, and placed in the upper right corner of their corresponding frames for display. As observed, our tracker can successfully tracks the targets in all challenging examples, featuring camera motion (i.e, in all sequences), occlusion (i.e, uav0000164_00000_s), and background clutter (i.e, S1606). Our method not only exhibits superior performance but also delivers more visually pleasing results in these scenarios, further validating the effectiveness of the proposed method for UAV tracking.
Table 4: Ablation study on the impact of MBRV and DEEM on the baseline tracker’s performance.
Tracker MBVR DEEM UAVDT VisDrone2018 UAV123 Avg Avg.FPS
Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ.
BDTrack-DeiT 78.6 56.7 81.6 62.2 83.7 66.1 81.3 61.7 228.7
$\checkmark$
84.5
61.4
86.1
65.1
85.3
67.0
85.3
64.5
$\checkmark$ $\checkmark$ 84.1 61.0 85.2 64.3 84.8 66.7 84.7 64.0
279.4
↑22.3%
4.7 Ablation Study
@@
4.7.1 Impact of learning blur robust ViTs with dynamic early exit
@@
We add the proposed components MBRV and DEEM into the baselines one by one to evaluate their effectiveness. The evaluation results on three benchmarks are shown in Table 4. As GPU speed of the baseline and its enhanced version by MBRV are theoretically equal, we provide only the speed of the baseline to eliminate potential nuances arising from randomness. As can be seen, the inclusion of MBRV leads to significant improvements in both Prec. and Succ. for the baseline tracker. Specifically, the inclusion of the MBRV leads to a significant increase of over 4.0% in average Prec. and 2.8% in average Succ. for the baseline. The improvement in performance after adding the MSE loss can be attributed to the enhanced feature consistency between clear and blurred templates. By enforcing feature similarity, the model learns more robust representations that generalize well to different blur conditions and focuses on discriminative features, leading to overall better performance beyond just handling blurred scenes. When DEEM is further integrated, consistent improvements in GPU speeds are observed, accompanied by only slight decreases in Prec. and Succ. Specifically, the GPU speeds of all baselines see increases of above 22.0%. Despite the marginal drop in Prec. and Succ. (the average drops are less than 0.6%), the overall pattern of enhanced GPU speeds and minor sacrifices in tracking performance validates the effectiveness of DEEM in optimizing tracking efficiency.
@@
Table 5: Ablation study on the weighting of loss $\mathcal{L}_{br}$ in training BDTrack-DeiT by varying $\rho$ from $0.5\times 10^{-4}$ to $5.0\times 10^{-4}$.
UAVDT VisDrone2018 UAV123 UAV123@10fps Avg.
$\rho\times 10^{-4}$ Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ.
0.5 83.4 60.0 83.7 63.5 83.1 65.4 83.2 65.6 83.3 63.6
0.6 82.3 59.3 84.5 63.9 83.7 66.0 83.0 65.4 83.4 63.6
0.7 83.0 59.6 85.8 64.6 84.1 66.5 83.3 66.2 84.0 64.2
0.8 83.2 59.8 81.1 61.9 83.5 65.8 82.3 64.9 82.5 63.1
0.9 82.5 59.7 84.6 63.9 84.0 66.1 83.1 65.6 83.6 63.8
1.0 84.1 61.0 85.2 64.3 84.8 66.7 83.5 65.9 84.4 64.5
2.0 80.4 58.2 86.1 64.8 84.0 65.9 84.9 66.5 83.8 63.8
3.0 83.8 60.7 85.6 64.5 82.4 64.7 80.2 63.6 83.0 63.3
4.0 82.4 59.4 84.5 63.8 84.5 66.5 80.0 63.5 82.8 63.3
5.0 82.6 59.7 83.9 63.6 84.0 66.0 82.3 65.1 83.2 63.6
@@
4.7.2 Study on weighting the loss for learning motion blur robust ViTs (MBRV)